OpenStudy (anonymous):

if f(x)= sqrt x^2-1 and g(x) = sqrt x-1, which expression represent f(x)/g(x) for x>1

OpenStudy (anonymous):

the one that looks like \[\frac{f(x)}{g(x)}\] with \(f(x)=\sqrt{x^2-1}\) and \(g(x)\) replaced by \(\sqrt{x-1}\)

OpenStudy (anonymous):

\[f(x)=\sqrt{x ^{2}-1} and g(x)\sqrt{x-1}, which expression represent \frac{ f(x) }{ g(x) } for x>1\]